OpenStudy (anonymous):

For what rational number "c" do the equations x^3+cx^2+3=0 and x^2+cx+1 = 0 have a common solution?

OpenStudy (mertsj):

c=-3

OpenStudy (anonymous):

can you explain how do you get it?

OpenStudy (mertsj):

\[x^3+cx^2+3=x^2+cx+1\] \[x^3+cx^2-x^2-cx+2=0\] \[x^3+(c-1)x^2-cx+2=0\]

OpenStudy (mertsj):